# Unscrambling of single-particle wave functions in systems localized through disorder and monitoring

### v1

@@ V Unscrambling without particle number conservation
of size $2L\times L$, where each column defines the coefficients of a Bogoliubov fermion. Similarly to the particle-conserving problem, we can define the inverse participation ratio of a particle $n$ as
@@
V.1 Disordered monitored free fermions without particle number conservation
We choose the model of monitored fermions without particle number conservation similar to those in Refs. [41, 42, 43, 44] but with an additional disordered field. Similarly to Sec. IV, this allows us to test the viability of the method not only in monitored quantum circuits, but also in the presence of the disorder. Comments parallel to Sec. IV will be applicable here – we expect that for low values of both measurement strength and the disorder strength, the model will show a critical entanglement behavior, while in other parts of the phase diagram, the area law will be present. The model in question is governed by the modified Anderson Hamiltonian,
@@ V.1 Disordered monitored free fermions without particle number conservation
The stochastic Schrödinger equation from Eq. (28) is now equivalent to the following evolution of the matrix $\mathbb{U}$:
@@
where the elements of $\mathbb{H}$ are defined through $H=\Psi^{\dagger}\mathbb{H}\Psi$ with the Nambu spinor $\Psi=(c_{1},...,c_{L},c^{\dagger}_{1},...,c^{\dagger}_{L})^{T}$; similarly for the elements of $\mathbb{M}$. Note the appearance of the factors of 2 in comparison to Eq. (29[61]. We initialize the wave function in the vacuum state $|0\rangle$, which is equivalent to $U=\mathbb{1}$ and $V=\mathbb{0}$, and evolve the state until the steady state is reached, with $t_{\text{fin}}=16L$ when $W\leq 1$ and $t_{\text{fin}}=16LW$ for $W>1$. The anisotropy is set to $\kappa=0.7$. The matrix $\mathbb{U}$ is subsequently unscrambled using the cost function of Eq. (35).
@@
We present the unscrambled single-particle wave functions for particles $\overline{\overline{|\psi^{\bullet}_{i}|^{2}}}$ (columns of matrix $U$) for $L=200$ in Fig. 8, where panel (a) shows example parameters for which conformal symmetry is present and panel (b) showcases the area-law phase. The corresponding wave functions for holes $\overline{\overline{|\psi^{\circ}_{i}|^{2}}}$ (columns of matrix $V$) are nearly identical to those of particles, and are therefore not shown in the figures. Surprisingly, the results do not differ significantly from our observations in the particle-conserving case. Specifically, we observe both conformally-symmetric delocalized wave functions in Fig. 8(a) (with clear signatures of $\overline{\overline{|\psi^{\bullet}_{i}|^{2}}}\sim\csc^{2}(i\pi/L)$ behavior), and exponentially localized orbitals in Fig. 8(b). Our findings strongly imply that even without particle number conservation, disordered monitored free fermions continue to exhibit a measurement-induced phase transition between an area law and a critical phase with logarithmic entanglement. Indeed, we anticipate that the resulting phase diagram will closely resemble that reported in Ref. [51]. We have approximated how such a phase diagram might appear in Fig. 9, where the unscrambled orbitals for $L=200$ were assessed to either fit conformal or localized behavior (see Appendix B for details). It is important to stress that a rigorous finite-size scaling analysis is necessary to pinpoint the exact location of the transition line. However, this crude estimation emphasizes two important features of this model: the existence of an entanglement transition and the non-monotonicity of the transition line. In previous literature [63, 51] this non-monotonic behavior has been associated with a mechanism, where a small amount of disorder could facilitate entanglement spreading, stabilizing the delocalized phase. Our results thus suggest that this mechanism is still present in the model without the $U(1)$ symmetry.
Figure 9: Phase diagram of the disordered monitored free fermions without particle number conservation. The crude estimate of the transition line (green) is done by estimating whether the behavior of the unscrambled single-particle wave functions is localized (blue circles) or conformal (orange squares). Undecided behavior is designated by green diamonds.
@@
VI Discussion and outlook
In this work, we have delved into the problem of single-particle representations of free-fermion wave functions. Specifically, our focus was revealing the underlying physics of free-fermion systems with localization. We have developed a method to find a representation in the most localized basis, where the proposed cost function is always decreasing and has a physical interpretation of maximizing the inverse participation ratio of each single-particle orbital. The method works well on paradigmatic examples of 1) conformally symmetric, 2) exponentially localized, and 3) power-law localized systems, and can be extended to encompass systems without particle number conservation. This methodology was shown to be important in understanding the delocalization-localization transition properties and give insights into the internal structure of free-fermion wave functions. Specifically, the method informs about the properties of the entanglement transition in monitored free-fermion models, and in models with a disorder, where the system exhibits Anderson localization.
Several interesting avenues emerge for further exploration of this work. Firstly, while we have successfully identified clear signatures deep within distinct phases, pinpointing the phase transition line based solely on single-particle wave functions remains a challenge. Therefore, the need for proper finite-size scaling analysis arises. It would be fascinating to investigate whether such an analysis can unveil universal properties of the transition, such as critical exponents. Secondly, the method should be readily extendable to free fermions in two and higher dimensions, where monitoring also forces a phase transition, albeit of a different nature [39, 55, 56]. Thirdly, the question remains whether it is possible to expand the unscrambling methodology to systems beyond non-interacting fermions. Bethe-ansatz solvable integrable models, characterized by states in the Slater determinant form, present an obvious choice for such exploration. Additionally, while most states cannot be precisely expressed as Slater determinants, many can be approximated as such. It would be intriguing to explore the implications of the unscrambling process for these approximations, and whether this would provide us with any further physical insights. As we venture into the topic of the interacting many-body models, the need for such insights becomes increasingly pronounced, together with the need for methods providing us with unique lenses – allowing us to unravel the intricate dynamics of complex quantum systems.
@@ Acknowledgements.
The author would like to thank Christopher J. Turner for useful discussions. M.S. was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 853368). The author acknowledges the use of the UCL Myriad High Performance Computing Facility (Myriad@UCL), and associated support services, in the completion of this work.
@@ Appendix A Unscrambling method and local minima of the multi-particle cost function
Figure 10: Difference $\varepsilon$ of cost functions between transformed single-particle orbitals $U^{\prime}=\exp(i\beta G)$ and the original $U$. (a) Example of the unscrambling when the result has not converged. $\varepsilon$ may be negative for some random matrices $G$ for low enough $\beta$, signifying that the result is not a minimum. (b) Converged result is numerically a local minimum, with $\varepsilon$ always positive. All panels use results for the disordered monitored free fermions without particle number conservation (see Sec. V) for $\gamma=0.05,W=4.5$, and show 100 random realizations of $G$.
@@
where $G$ is a Hermitian random matrix from the Gaussian unitary ensemble, and $\beta$ is a parameter controlling the closeness of $Q$ to the identity matrix. Applying $Q$ to the matrix of single-particle wave functions $U$ takes us away from the unscrambling result in a random direction, $U^{\prime}=UQ$. We then compare the difference $\varepsilon$ between the cost function for $U^{\prime}$ and that for $U$. If this difference is positive for any choice of $G$ and $\beta$, it confirms that the unscrambling outcome corresponds to a local minimum.
@@
Appendix B Phase diagram of the disordered monitored free fermions without particle number conservation
Figure 11: Density plot of the function $g$ that measures the difference between the results and the conformal behavior for the disordered monitored free fermions without particle number conservation.
@@ Appendix B Phase diagram of the disordered monitored free fermions without particle number conservation
where $\mathcal{L}$ encompasses the middle 80% of the sites. The result is shown in Fig. 11. The points in the red region are chosen as being close to critical, the green region is localized, and the blue region points are designated as intermediate. The transition line is then estimated as the boundary between the red and green regions.

### v3

@@ VI Unscrambling without particle number conservation
of size $2L\times L$, where each column defines the coefficients of a Bogoliubov fermion. Similarly to the particle-conserving problem, we can define the inverse participation ratio of an orbital $n$ as
@@
VI.1 Disordered monitored free fermions without particle number conservation
We choose the model of monitored fermions without particle number conservation similar to those in Refs. [65, 66, 67, 68] but with an additional disordered field. Similarly to Sec. V, this allows us to test the viability of the method not only in monitored quantum circuits, but also in the presence of the disorder. Comments parallel to Sec. V will be applicable here – we expect that for low values of both measurement strength and the disorder strength, the model will show a critical entanglement behavior, while in other parts of the phase diagram, the area law will be present. The model in question is governed by the modified Anderson Hamiltonian,
@@ VI.1 Disordered monitored free fermions without particle number conservation
The stochastic Schrödinger equation from Eq. (28) is now equivalent to the following evolution:
@@
where the elements of $\mathbb{H}$ are defined through $H=\Psi^{\dagger}\mathbb{H}\Psi$ with the Nambu spinor $\Psi=(c_{1},...,c_{L},c^{\dagger}_{1},...,c^{\dagger}_{L})^{T}$; similarly for the elements of $\mathbb{M}$. Note the appearance of the factors of 2 in comparison to Eq. (29[85]. We initialize the system in the vacuum state $|0\rangle$, which is equivalent to $U=\mathbb{1}$ and $V=\mathbb{0}$, and evolve the state until the steady state is reached, with $t_{\text{fin}}=16L$ when $W\leq 1$ and $t_{\text{fin}}=16LW$ for $W>1$. The anisotropy is set to $\kappa=0.7$. The matrix $\mathbb{U}$ is subsequently unscrambled using the cost function of Eq. (35).
@@
We present the unscrambled single-particle wave functions for particles $\overline{\overline{|\psi^{\bullet}_{i}|^{2}}}$ (columns of matrix $U$) for $L=200$ in Fig. 9, where panel (a) shows example parameters for which conformal symmetry is present and panel (b) showcases the area-law phase. The corresponding wave functions for holes $\overline{\overline{|\psi^{\circ}_{i}|^{2}}}$ (columns of matrix $V$) are nearly identical to those of particles, and are therefore not shown in the figures. Surprisingly, the results do not differ significantly from our observations in the particle-conserving case. Specifically, we observe both conformally-symmetric wave functions in Fig. 9(a) (with clear signatures of $\overline{\overline{|\psi^{\bullet}_{i}|^{2}}}\sim\csc^{2}(i\pi/L)$ behavior), and exponentially localized orbitals in Fig. 9(b). Our findings strongly imply that even without particle number conservation, disordered monitored free fermions continue to exhibit a measurement-induced phase transition between an area law and a critical phase with logarithmic entanglement. Indeed, we anticipate that the resulting phase diagram will closely resemble that reported in Ref. [75]. We have approximated how such a phase diagram might appear in Fig. 10, where the unscrambled orbitals for $L=200$ were assessed to either fit conformal or localized behavior (see Appendix C for details). It is important to stress that a rigorous finite-size scaling analysis is necessary to pinpoint the exact location of the transition line. However, this crude estimation emphasizes two important features of this model: the existence of an entanglement transition and the non-monotonicity of the transition line. In previous literature [87, 75] this non-monotonic behavior has been associated with a mechanism, where a small amount of disorder could facilitate entanglement spreading, stabilizing the critical phase. Our results thus suggest that this mechanism is still present in the model without the $U(1)$ symmetry.
Figure 10: Phase diagram of the disordered monitored free fermions without particle number conservation. The estimate of the transition line (green) is done by estimating whether the behavior of the unscrambled single-particle wave functions is localized (blue circles) or conformal (orange squares). Undecided behavior is designated by green diamonds.
@@
VII Discussion and outlook
In this work, we have investigated the problem of single-particle representations of free-fermion wave functions. Specifically, our focus was revealing the underlying physics of free-fermion systems with localization. We have developed a method to find a representation in the most localized basis, where the proposed cost function is always decreasing and has a physical interpretation of maximizing the inverse participation ratio of each single-particle orbital. The method works well on paradigmatic examples of 1) conformally symmetric, 2) exponentially localized, and 3) power-law localized systems, and can be extended to encompass systems without particle number conservation. This methodology was shown to be important in understanding the localization transition properties and give insights into the internal structure of free-fermion wave functions. Specifically, the method informs about the properties of the entanglement transition in monitored free-fermion models, and in models with a disorder, where the system exhibits Anderson localization.
Several interesting avenues emerge for further exploration of this work. Firstly, while we have successfully identified clear signatures deep within distinct phases, pinpointing the phase transition line based solely on single-particle wave functions remains a challenge. Therefore, the need for proper finite-size scaling analysis arises. The scaling could, for example, be performed using the IPR of the unscrambled orbitals, or some other associated property. It would be fascinating to investigate whether such an analysis can unveil universal properties of the transition, such as critical exponents. Secondly, the method is readily extendable to free fermions in two and higher dimensions, where monitoring also forces a phase transition, albeit of a different nature [63, 79, 80].
Thirdly, the question remains whether it is possible to expand the unscrambling methodology to systems beyond non-interacting fermions. Bethe-ansatz solvable integrable models, characterized by states in the Slater determinant form (such as in Ref. [88]), present an obvious choice for such exploration. Furthermore, while most states cannot be precisely expressed as Slater determinants, many can be approximated as such. It would be intriguing to explore the implications of the unscrambling process for these approximations, and whether this would provide us with any further physical insights. In the context of many-body quantum systems, the one-particle density matrix (OPDM) description [17, 89, 90, 91] has proven useful for understanding many-body localization. Perhaps the unscrambling method could be applied to the natural orbitals (OPDM eigenvectors), although this may necessitate adapting the cost function to incorporate the information about orbital occupations (OPDM eigenvalues). Moreover, the interpretation of this procedure would require further consideration, albeit naively the unscrambling should yield the “maximally” localized set of orbitals. As we venture into the topic of interacting many-body models, the need for physical insights becomes increasingly pronounced, together with the need for novel methods that allow us to unravel the intricate dynamics of complex quantum systems.
All relevant data present in this publication can be accessed at [92].
@@ Acknowledgements.
The author would like to thank Christopher J. Turner for useful discussions. M.S. was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 853368). The author acknowledges the use of the UCL High Performance Computing Facilities (Myriad@UCL and Kathleen@UCL), and associated support services, in the completion of this work.
@@ Appendix A Unscrambling method and local minima of the multi-particle cost function
where $G$ is a Hermitian random matrix from the Gaussian unitary ensemble, and $\beta$ is a parameter controlling the closeness of $Q$ to the identity matrix. Applying $Q$ to the matrix of single-particle wave functions $U$ takes us away from the unscrambling result in a random direction, $U^{\prime}=UQ$. We then compare the difference $\varepsilon$ between the cost function for $U^{\prime}$ and that for $U$. If this difference is always positive for any choice of $G$ and $\beta$, it confirms that the unscrambling outcome corresponds to a local minimum.
@@
Figure 11: Difference $\varepsilon$ of cost functions between transformed single-particle orbitals $U^{\prime}=\exp(i\beta G)$ and the original $U$. (a) Example of the unscrambling when the result has not converged. $\varepsilon$ may be negative for some random matrices $G$ for low enough $\beta$, signifying that the result is not a minimum. (b) Converged result is numerically a local minimum, with $\varepsilon$ always positive. All panels use results for the disordered monitored free fermions without particle number conservation (see Sec. VI) for $\gamma=0.05,W=4.5$, and show 100 random realizations of $G$.
Appendix B Single-particle eigenvectors of the Anderson model
Figure 12: Unscrambled orbitals (solid lines) of the one-dimensional Anderson model on a periodic chain of size $L=204$, compared with the averaged single-particle eigenfunctions in the middle 10% (dashed lines) and edge 5% (dotted lines) of the spectrum.
In this Appendix, we compare the behavior of single-particle eigenvectors of the 1D Anderson model with the unscrambled orbitals of Sec. III.1. The eigenvectors can be obtained by diagonalizing the Hamiltonian matrix $\mathbb{H}$ of size $L\times L$, defined below Eq. (19). Conceptually, the evolution of the orbitals from Sec. III.1 explores the bulk of the Hilbert space, hence we expect the behavior to be representative of the system near $E\approx 0$ (middle of the spectrum). However, the single-particle eigenvectors of the Anderson Hamiltonian are quite special points of the Hilbert space, hence it is hard to make a direct comparison. Fig. 12 shows that the single-particle eigenfunctions generically decay more rapidly than the unscrambled orbitals. Single-particle eigenvectors at the edge of the spectrum exhibit a more rapid decay than those near the middle of the spectrum ($E\approx 0$).
Appendix C Phase diagram of the disordered monitored free fermions without particle number conservation
Figure 13: Density plot of the function $g$ that measures the difference between the results and the conformal behavior for the disordered monitored free fermions without particle number conservation. The transition line (in white) estimates the transition point between the conformal behavior and the area law.
@@ Appendix C Phase diagram of the disordered monitored free fermions without particle number conservation
where $\mathcal{L}$ encompasses the middle 80% of the sites. We then perform this calculation for a grid $W\in$0.0, 0.5, 1.5, 3.0, 4.5, 6.0, 7.5, 9.0}, $\gamma\in$0.05, 0.5, 1.5, 2.0, 3.0, 4.0} at the system size of $L=200$. We do not consider the line $\gamma=0$ corresponding to an Anderson-localized model, since it most likely follows a separate behavior mimicking Ref. [75], where there is an abrupt change as soon as any monitoring is introduced.
The result is shown in Fig. 13. The points in the red region are chosen as being close to critical, the green region is localized, and the blue region points are designated as intermediate. The transition line is then estimated as the boundary between the red and green regions.